# Knowledge Augmentation in Federation: Rethinking What Collaborative Learning Can Bring Back to Decentralized Data

### v2

@@ Abstract
Data, as an observable form of knowledge, has become one of the most important factors of production for the development of Artificial Intelligence (AI). Meanwhile, increasing legislation and regulations on private and proprietary information results in scattered data sources also known as the “data islands”. Although some collaborative learning paradigms such as Federated Learning (FL) can enable privacy-preserving training over decentralized data, they have inherent deficiencies in fairness, costs and reproducibility because of being learning-centric, which greatly limits the way how participants cooperate with each other. In light of this, we present a knowledge-centric paradigm termed Knowledge Augmentation in Federation (KAF), with focus on how to enhance local knowledge through collaborative effort. We provide the suggested system architecture, formulate the prototypical optimization objective, and review emerging studies that employ methodologies suitable for KAF. On our roadmap, with a three-way categorization we describe the methods for knowledge expansion, knowledge filtering, and label and feature space correction in the federation. Further, we highlight several challenges and open questions that deserve more attention from the community. With our investigation, we intend to offer new insights for what collaborative learning can bring back to decentralized data.
@@
Decentralized data, knowledge extraction, knowledge augmentation, federated learning, data representation, data engineering.
@@ I Introduction
As Artificial Intelligence (AI) exhibits its potential in transforming human society, we are entering a new era where the amount of data and the value generated from data both increase in an unprecedented speed. Statistics show that the total volume of data or information created, captured, copied and consumed globally has experienced a tenfold growth in the past decade and is expected to reach 394 zetabytes in 2028 [1]. This is in large part attributed to the development of data acquisition and sensing techniques as well as the pervasive deployment of Machine Learning (ML) models. For example, research predicts that a single autonomous vehicle can generate as much as 10 gigabytes data per second in the future [2]. The ever-increasing availability of data makes it possible for us to build and train more advanced models such as the Broad Learning System (BLS) [3, 4] and Large Language Models (LLMs) [5, 6].
Figure 1: The conceptual overview of Knowledge Augmentation in Federation (KAF) as a knowledge-centric paradigm for local domain enhancement.
While AI applications are more data-hungry than ever, privacy and proprietary issues arise. The accessibility of private data is in many cases limited by legislation and regulations such as the General Data Protection Regulation (GDPR) [^1] https://gdpr-info.eu/ and the California Consumer Privacy Act (CCPA) [^2] https://www.oag.ca.gov/privacy/ccpa. This to a great extent leads to a paradigm shift from centralized learning to collaborative learning from decentralized data. To address the problem of “data islands”, McMahan et al. [7] from Google proposed Federated learning (FL) [8], a privacy-preserving paradigm based on the concept of Federated Optimization designed for collaborative learning over autonomic user devices. The general workflow of FL involves global model distribution, local training and model upload, and global model aggregation. Essentially, the key factors that distinguish Federated learning (or Federated Optimization) concern the massive system scale, restricted communication and most importantly, the non-Independent and Identically Distributed (non-IID) data assumption [9]. Although extensive research has been conducted to enhance the efficacy and efficiency of FL in diverse scenarios, existing methods hardly go beyond model-based knowledge sharing where models [^3] Models and model parameters are used interchangeably if no ambiguity is caused., model updates, or gradients are exchanged between different parties as the carrier of knowledge. This helps in protecting the raw data but also results in several deficiencies:
@@
  • Prohibitive costs: Models are “heavy”, and so are the model updates and the gradients. They are generally in the same tensor structure and can take up to gigabytes of network traffic for each transmission. Besides, the cost for local training cannot be overlooked considering the problem of slow convergence over non-IID data [10].
@@
  • Poor reproducibility: models are the carriers of knowledge but are hard to reproduce in massive collaboration due to the randomness in participation. Any participant has to request massive cooperation in order to reproduce a model from scratch.
Is it possible to find an alternative to model-centric collaboration and tackle the above-mentioned issues? In this paper we offer our answer by presenting a knowledge-centric collaboration paradigm termed Knowledge Augmentation in Federation (KAF) or Federated Knowledge Augmentation. The key idea behind KAF is to redefine the way that participants collaborate in a federated system by shifting the focus from learning models to enhancing knowledge across the domains, as illustrated in Fig. 1. Technically, the key difference that KAF brings to the field mainly concerns i) the introduction of knowledge-centric objectives for collaborative optimization in the federated system, and ii) the adoption of more flexible carriers of knowledge without compromising privacy or proprietary rights. More details will be elaborated in Sec. III.
@@
On the one hand, typical distributed learning paradigms seek to build a strong model by aggregating models, gradients or other forms of knowledge from separate domains. Even in heterogeneity-aware approaches such as personalized Federated Learning (pFL) [11], the optimization is still mostly limited to task-specific models which are hardly generalizable. By contrast, Knowledge Augmentation in Federation emphasizes the enhancement of local knowledge associated with each individual participant. On the other hand, The knowledge brought in can take whatever form, e.g, models, gradients, predictions, synthetic data, etc., as long as it serves to enrich or refine the knowledge implied in local data (Fig. 2). This calls for new protocols that enable cost-effective exchange of knowledge. We will elaborate on these points in Sec. III. Compared to existing paradigms, KAF has potential advantages summarized as follows:
  1. 1.
    Improving global knowledge coherence across non-IID data holders,
  2. 2.
    Enhancing local knowledge representations in quantity, quality and utility for better local training or independent learning, and
  3. 3.
    Knowledge-centric collaboration for augmenting local domains with learning-friendly data rather than task-specific models.
Figure 2: The conceptual relationship between data, models and knowledge.
@@
Moreover, we believe knowledge augmentation in federation shows a new roadmap for autonomic intelligence, where each participant can learn independently from knowledge gained collaboratively. In fact, there exist a number of studies that adopted inspiring methodologies in relevance to KAF, but a comprehensive review is so far missing in the literature. As far as we know, this review is the first to assemble related studies in the field under the concept of KAF.
I-A Contributions and Related Surveys
Extensive research has been carried out in the fields of deep learning (DL), representation learning, and data engineering. Meanwhile, demands for data privacy have given rise to broad interest in collaborative learning paradigms such as federated learning, swarm learning and fog learning. In the following we briefly review the representative surveys under these topics and then highlight the difference made in our work.
Deep learning and representation learning. There exist a number of comprehensive surveys on DL-related techniques and applications. Dong et al. [12] briefly reviewed the development of Deep Neural Networks (DNNs) as well as other fundamental elements for DL such as activation functions, model architectures and programming libraries. Similarly, Sarker [13] categorized typical DL approaches in a taxonomy based on the availability of labels. Realistic tasks for supervised learning, unsupervised learning, hybrid learning and other learning paradigms are systematically reviewed. From a different angle, Alzubaidi et al. [14] paid attention to the impact of data scarcity on DL model training and how to deal with it. Besides, how to learn strong representations from the data is believed to be the key to DL algorithm design [15]. Considering the connection between data representations with domain-specific knowledge, surveys have been conducted with focuses on various domains such as network representation learning [16], multi-modal representation learning [17], and graph representation learning [18].
Data engineering and management. The development of Big Data techniques not only paves the way for complex learning systems, but also calls for more learning-centric data engineering. Generally speaking, data engineering entails the whole lifecycle of data management from acquisition and pre-processing to utilization, each with significant impact on the outcome of learning [19]. Among the problems in the field, data imbalance has attracted the most attention, and a diversity of pre-processing and learning-based methods were investigated in the literature [20, 21]. In addition, surveys such as [22, 23, 24] shed light on the bias and fairness in data. However, collaborative engineering of knowledge beyond data is yet to be covered in existing surveys.
Federated learning and its variants. Kairouz et al. [9] presented one of most comprehensive surveys on FL, covering the efficiency, effectiveness, privacy, robustness and fairness issues. They also highlighted the challenges related to ill-distributed data for both cross-device and cross-silo FL scenarios. Some surveys focus on particular aspects of an FL system such as privacy preservation [25], decentralization [26], and heterogeneous FL [27]. Application-specific studies are also reviewed in the literature, e.g., FL for Internet of Things (IoT) [28], smart cities [29], and healthcare [30, 31, 32]. Ji et al. [33] termed the variants of FL paradigms as “Federated X Learning”, which represents the fusion of FL with other emerging ML paradigms. Nonetheless, these surveys, without any exception, are learning-centric and did not investigate what FL can bring back to the data.
Table I highlights the aspects that distinguish this review paper from the rest in the literature.
@@ I-A Contributions and Related Surveys
Knowledge-centric collaboration learning data & feature engineering
DL & representation learning (e.g., [12, 13, 14, 15, 16, 17, 18])
Data engineering & management (e.g., [19, 20, 21, 22, 23, 24])
Federated learning (e.g., [9, 25, 26, 27, 28, 29, 30, 31, 32, 33])
@@
Our contributions. In this paper we advocate a knowledge-centric vision and focus on what learning can bring back to data in a federated system. Main contributions of this work are summarized as follows:
  • A revisit to data processing and learning across “data islands” from the perspective of knowledge.
  • New understanding of collaborative learning as a bridge to knowledge augmentation in federation.
  • Insights for what KAF can bring up for the evolution of AI and the new regimes of collaborative intelligence.
II Background and Preliminaries
In this section we provide a brief introduction to the basics of federated learning and data engineering, which helps to better understand the ideas and topics covered in our work.
II-A Federated Learning
Federated learning is a distributed machine learning paradigm for privacy-preserving collaboration across a multitude of devices. The basic system model of FL includes user devices (often termed clients) that possess private data and a central server that coordinates the learning process [34]. The standard process of FL [7] is organized in rounds each comprised of the following steps: i) Client selection: a fraction of clients are selected for the upcoming round of training; ii) Model distribution: the server synchronizes with the selected clients by sending them the global model (consensus model); iii) Local training: each selected client trains the model received on its local data using gradient-based methods, which produces a local model; iv) Model upload: selected clients upload their local models to the server; and v) Model aggregation: the server aggregates the local models into a new global model. These steps cycle for a number of rounds until some criterion is met.
A unique challenge that distinguishes federated learning from other learning paradigms is the multi-party ownership of data. This significantly complicates the way we operate across multiple disjoint sets of data, i.e., “data islands”. On the one hand, in the FL setting we need to deal with a second layer of data distribution, that is, a distribution of similar but different domains represented by all the client-owned local datasets in the system. On the other hand, we still face the notorious problems, such as data imbalance, sample corruption and label noise, probably within each of the local datasets. More importantly, the orchestrator (e.g., a central server in most cast) is typically agnostic of local data distribution due to privacy constraints.
II-B Data Engineering
Data engineering in a general sense refers to the practice in the management of data spanning its whole lifecycle [23]. Data engineering mainly deals with the problems in data collection and integration, data storage, data processing, and data analytics [^4] Systematic aspects such as data pipeline automation are also elements of the theme but do not fall within the scope of this paper.. The boom of ML and AI applications has brought increasing attention to learning-centric data engineering methods particularly including data augmentation, data balancing and feature engineering. In the following we briefly explain these concepts.
Data augmentation and balancing. The goal of data augmentation is to facilitate learning or analytics by artificially enhancing a dataset in size and sample diversity. Widely adopted in image, text and rational data pre-processing, data augmentation usually involves the generation of synthetic samples and the transformation of raw data. On the one hand, data synthesis introduces artificial samples to the dataset by exploiting interpolation, extrapolation, statistical models, generative models, or any combination of these techniques [35]. Note that data synthesis is also a common strategy for data balancing or re-sampling where the distribution of samples is amended by over-sampling the minority class. This is the very rationale behind the classic methods such as Synthetic Minority Oversampling Technique (SMOTE) [36], ADAptive SYNthetic (ADASYN) [37], and Mixup [38]. On the other hand, data transformation helps to diversify the features of data by applying certain types of transforms. The transformation of data are usually domain-specific, e.g., rotation, flipping and rescaling for image data, and random deletion, insertion and rephrasing for text data.
Feature Engineering (FE). Raw data are more often than not represented by attributes that can be noisy, highly correlated and even corrupted. The goal of feature engineering is to create a new set of features that better represent the knowledge implied and filter out the noise and misinformation. Typical recipes for feature engineering includes feature selection and feature transformation and extraction. Specifically, feature selection reduces the dimensionality of the feature space by filtering for the most important features, which is usually performed based on statistical evidence or measures from information theory [39]. As for feature transformation and extraction, it not only applies numerical operation(s) like rescaling to individual features, but also involves re-encoding the data with new combinations of features [40]. Considering the target value (label) as a special feature, methods such as label correction and generation can also be categorized as the implementation of FE. In practice, domain knowledge can be leveraged in the FE process. The confluence of feature engineering and representation learning unleashes the power of AI over massive data.
@@ III Conceptual Overview
In this section we provide a holistic view of the proposed KAF paradigm. We start with introducing the (suggested) system architecture and move on to the problem setting. This is followed by the outline of key approaches that will be elaborated in Secs. IV, V and VI.
@@ III-A System Architecture
As shown in Fig. 1, the system model of KAF is similar to that of (horizontal) federated learning. To be specific, a central server is deployed as the coordinator at the center of the hub-and-spoke topology. A multitude of participants $\mathcal{K}=\{u_{1},u_{2},\ldots\}$ connect to the server to send and receive shared knowledge. Each participant $u_{k}$ has the exclusive access to its local data $D_{k}$ that is not allowed to be shared for privacy or proprietary reasons.
A key difference is that each participant in KAF is not necessarily capable of training deep ML models. This is because the optimization objective for KAF is not learning-centric but knowledge-centric, which will be elaborated in the following section.
III-B Problem Statement
The fundamental problem behind privacy-sensitive collaboration is the “data islands” that causes non-IID data across the participants (i.e., clients in FL). From the view of knowledge, this problem can be interpreted as a global domain segmented into multiple overlapped sub-domains [^5] Also simply referred to as (local) domains if no ambiguity is caused. corresponding to the local data partitions on the clients.
The proposed paradigm, namely knowledge augmentation in federation, seeks to address the problem by exchanging and enhancing the knowledge in private data independent on task-specific model learning. In other words, the primary goal of KAF is to maximize the effective capacity of knowledge for each participant [^6] The terms participant and client are used interchangeably. $u_{k}\in\mathcal{K}$ that holds a private dataset $D_{k}$, where $\mathcal{K}$ represents the entire population in the system. In addition, we also expect each local domain to align with the global domain through collaboration but without deviating too much from its original data distribution $\mathcal{D}_{k}$. This translates to a prototypical objective function expressed in (1),
$\max_{\pi_{k}}F_{k}=\underbrace{\nu\big{(}\Theta_{\pi_{k}}(D_{k};\mathcal{K})\big{)}}_{enhancement}-\\ \alpha_{k}\Big{(}\beta_{k}\cdot\underbrace{\sum_{u_{i}\in\mathcal{K}/u_{k}}d\big{(}\Theta_{\pi_{k}}(D_{k};\mathcal{K}),\Theta_{\pi_{k}}(D_{i};\mathcal{K})\big{)}}_{alignment\;loss}+\\ (1-\beta_{k})\cdot\underbrace{d\big{(}\Theta_{\pi_{k}}(D_{k};\mathcal{K}),D_{k}\big{)}}_{regularization}\Big{)},$ (1)
where the function $\nu(\cdot)$ evaluates the capacity of knowledge given a dataset, $\Theta_{\pi}(\cdot;\mathcal{K})$ denotes the operations following a strategy (or protocol) $\pi$ on local data collaboratively performed across the federation $\mathcal{K}$, and $d(\cdot,\cdot)$ measures the distributional distance (e.g., Kullback-Leibler (KL) divergence) between two datasets. The two client-specific hyper-parameters $\alpha_{k}$ and $\beta_{k}$ weight the three terms in the objective.
Collectively, one can derive a system-wide objective from (1) if all the participants are cooperative and stick to their strategies $\{\pi_{1},\pi_{2},\ldots\}$:
$\max_{\{\pi_{1},\pi_{2},\ldots\}}\sum_{k=1}^{|\mathcal{K}|}F_{k}$ (2)
Besides, objective (1) indicates that each participant in KAF is self-motivated and contributes towards its own benefit. Particularly, the augmented knowledge represented in local data makes it possible for each $u_{k}$ to carry out independent model training or any other downstream tasks, making the cooperation knowledge-centric rather than learning-centric. On this point, KAF also differs from the objective of pFL [41] where each client aims to minimize a task-specific loss w.r.t. a specific model.

### v3

# Task-Agnostic Federation over Decentralized Data: Research Landscape and Visions

@@ Abstract
Increasing legislation and regulations on private and proprietary information results in scattered data sources also known as the “data islands”. Although Federated Learning-based paradigms can enable privacy-preserving collaboration over decentralized data, they have inherent deficiencies in fairness, costs and reproducibility because of being learning-centric, which greatly limits the way how participants cooperate with each other. In light of this, we investigate the possibilities to shift from resource-intensive learning to task-agnostic collaboration especially when the participants have no interest in a common goal. We term this new scenario as Task-Agnostic Federation (TAF), and investigate several branches of research that serve as the technical building blocks. These techniques directly or indirectly embrace data-centric approaches that can operate independently of any learning task. In this article, we first describe the system architecture and problem setting for TAF. Then, we present a three-way roadmap and categorize recent studies in three directions: collaborative data expansion, collaborative data refinement, and collective data harmonization in the federation. Further, we highlight several challenges and open questions that deserve more attention from the community. With our investigation, we intend to offer new insights about how autonomic parties with varied motivation can cooperate over decentralized data beyond learning.
@@
Generic knowledge, Decentralized data, knowledge exchange, data engineering, federated systems.
@@ I Introduction
As Artificial Intelligence (AI) exhibits its potential in transforming human society, we are entering a new era where the amount of data and the value generated from data both increase in an unprecedented speed. Statistics show that the total volume of data or information created, captured, copied and consumed globally has experienced a tenfold growth in the past decade and is expected to reach 394 zetabytes in 2028 [statista2028]. A research predicted that a single autonomous vehicle can generate as much as 10 gigabytes data per second in the future [yang2020vehicle]. This signifies two trends in the development of Machine Learning (ML) and data mining techniques. On the one hand, vast data lays the foundation for large-scale learning systems [guo2024bls, guo2025deepseek, zhong2024evaluation]. On the other hand, knowledge sharing across multiple data sources matters more than ever [zhao2024domain, guan2025effective].
In reality, the accessibility of private data is in many cases limited by legislation and regulations such as the General Data Protection Regulation (GDPR) [^1] https://gdpr-info.eu/ and the California Consumer Privacy Act (CCPA) [^2] https://www.oag.ca.gov/privacy/ccpa. This to a great extent leads to a paradigm shift from centralized learning to collaborative training over decentralized data. The most popular solution to “data islands” is Federated learning (FL) [mcmahan2016arxiv, mcmahan2017communication], which was developed on the basis of Federated Optimization for collaborative model training [kairouz2021fl]. Although extensive researches have been conducted following the FL paradigm, they hardly go beyond learning-centric, task-specific knowledge sharing. This protects the raw data but results in several deficiencies:
@@
  • Prohibitive costs: Models are “heavy”, and so are the model updates and the gradients. They are generally in the same tensor structure and can take up to gigabytes of network traffic for each transmission. Besides, the cost for local training cannot be overlooked considering the problem of slow convergence over non-IID data [amiri2022conv].
@@
  • Poor reproducibility: models can carry knowledge but are hard to reproduce in massive collaboration due to the randomness in participation. Any participant has to request massive cooperation in order to reproduce a model from scratch.
Figure 1: The conceptual overview of Task-Agnostic Federation (TAF), which is a data-centric scenario that liberates the participants from any common goal and employs flexible protocols for generic knowledge exchange.
Is it possible to find an alternative to task-oriented collaboration and tackle the above-mentioned issues? In this paper we offer our answer by describing a new scenario termed Task-Agnostic Federation (TAF), under which we gather emerging studies that enable new ways of collaboration through the exchange of generic knowledge in data. TAF liberates the participants by allowing for diverse motivation of collaboration in a federated system. This is done by shifting the focus from learning models to enhancing data across the domains, as illustrated in Fig. 1. This generalizes the benefits of collaboration because the enhanced data can take various forms and serve for any downstream purposes beyond learning. The concept of TAF will be elaborated in Sec. III.
@@
Traditional distributed learning paradigms seek to build a strong model by aggregating models, gradients or other forms of knowledge from separate domains. Even in heterogeneity-aware approaches such as personalized Federated Learning (pFL) and Federated Multi-Task Learning (FMTL) [rahman2024tvt, liu2025pfed, skovajsova2025fmtl], the optimization is still mostly limited to task-specific models which are hardly generalizable. By contrast, Task-Agnostic Federation emphasizes the enhancement of private data with generic knowledge across the system. Particularly, The knowledge brought in can take whatever form, e.g, statistical information, synthetic data, generators, class prototypes, predictions, etc., as long as it serves to enrich or refine the knowledge implied in local data (Fig. 2). This calls for new protocols that enable cost-effective exchange of knowledge under privacy constraints. Compared to existing paradigms, TAF has advantages in several aspects. First, it directly enhances local data in quantity, quality or utility. Second, it brings generic, reusable knowledge to users that can be utilized later for diverse custom purposes. Besides, it leads to sustainable gain for every participating party despite how much they are engaged [^3] By contrast, FL relies on persistent client engagement to produce strong models.. This is practically important for the participants who have very limited resources to invest.
@@
In fact, an increasing number of studies have adopted inspiring methodologies in relevance to TAF, but a comprehensive review is so far missing in the literature. As far as we know, this review is the first to assemble related studies in the field under this concept.
Our contributions. In this paper we advocate a data-centric vision and focus on how multiple parties can cooperate without a common goal in a federated system. We believe TAF distinguishes itself as a new roadmap for autonomic intelligence, where each participant seeks and shares knowledge that is generic and reusable. The value of this article is summarized as follows:
  • A revisit to multi-party collaboration across “data islands” from the standpoint of task-irrelevant interest and long-lasting benefits.
  • New understanding of collaboration as a bridge to augmenting, refining and transforming local data for general purposes rather than learning.
  • Literature review and insights for what TAF can bring up for the new forms of federated systems where autonomy matters.
@@
Theme Surveys Task-agnostic Collaboration Learning Local benefit
Data engineering & knowledge management (e.g., [balayn2021biasunfairness, cui2025kadl, xie2024dataquality, gonzales2023synth, caton2024fairness, mehrabi2022biasfairness])
Federated learning and variants (e.g., [kairouz2021fl, chen2024flprivacy, martinez2023decenfl, sabah2024model, ye2023heterofl, ji2024fxl, skovajsova2025fmtl]) ✓(pFL&FMTL)
Federated analytics (e.g., [wang2025fa, wang2022fa, elkordy2023federated])
@@
Figure 2: Data is the raw material and the most flexible form of knowledge. Information presented in statistics, patterns and models is derived from data and carries a specific subset of knowledge.
II Background and Related Surveys
In this section we briefly review some closely related concepts and draw comparison with related surveys to highlight our unique perspectives.
Data engineering and knowledge management. Data engineering is something usually “behind the screen” but vital to any learning or mining pipeline [balayn2021biasunfairness], whilst the management of knowledge is pervasive in practice spanning from data abstraction to expertise transfer. The rise of these topics dates back to decades ago when rule-based symbolism and expert systems prevailed. An early work by Studer et al. [studer1998knowledge] signified a shift towards modeling view and outlined two diverged approaches to refining knowledge, whilst a recent survey by Cui et al. [cui2025kadl] reviews how deep learning can be augmented by domain knowledge. A diversity of data-centric studies were also investigated in the literature, with topics covering data balancing [xie2024dataquality, gonzales2023synth, chawla2002smote], feature engineering [chandrashekar2014survey, salau2019extract] as well as data bias and fairness issues [caton2024fairness, balayn2021biasunfairness, mehrabi2022biasfairness]. However, new scenarios like collaborative data engineering and decentralized knowledge augmentation is yet to be covered in existing surveys.
Federated learning and its variants. Federated learning is a distributed machine learning paradigm for privacy-preserving collaboration across a multitude of devices [bonawitz2019flsys]. The most unique challenge for FL is the multi-party ownership of data. Kairouz et al. [kairouz2021fl] presented one of the most comprehensive surveys on FL, covering the efficiency, effectiveness, privacy, robustness and fairness issues. Some surveys focus on particular aspects of an FL system such as privacy preservation [chen2024flprivacy], decentralization [martinez2023decenfl], personalized FL (pFL) [sabah2024model], and heterogeneous FL [ye2023heterofl]. Application-specific studies are also reviewed in the literature, e.g., FL for Internet of Things (IoT) [nguyen2021fliot], smart cities [pandya2023flcities], and healthcare [taha2023flhealthcare]. Ji et al. [ji2024fxl] termed the variants of FL paradigms as “Federated X Learning”, which represents the fusion of FL with other emerging ML paradigms. Nonetheless, these surveys, without any exception, overlooked what learning can bring back to the data.
Federated Analytics. Federated analytics (FA) [googleai2020federated], usually developed on information seeking, confidential computing and cryptography, aims to obtain global statistics with minimal private information leakage. The motivation and unique challenges for FA are discussed in [wang2022fa], while Wang et al. [wang2025fa] provide a comprehensive view of research landscape, with a taxonomy that entails analytical tasks, algorithms as well as privatization. Elkordy et al. [elkordy2023federated] review relevant techniques to FA from a query-oriented perspective. Yet, FA does not bring knowledge back to the data owners, making what we are interested in uncovered in these surveys.
In contrast to existing surveys, we advocate a task-agnostic vision and investigate how multiple parties in a federated system can benefit from data-centric cooperation. Table I highlights the aspects that distinguish this review paper from the rest in the literature.
@@ III Conceptual Overview
In this section we provide a holistic view of the proposed scenario of TAF, starting with a general system architecture and moving on to the problem setting. This is followed by the outline of key approaches that will be elaborated in Secs. IV, V and VI.
@@ III-A System Architecture
As shown in Fig. 1, the system model in TAF is similar to that of (horizontal) federated learning. Considering the diversity of data sources, a central server is usually needed for coordinating the scope and protocol of cooperation. On this hub-and-spoke topology, a multitude of participants $\mathcal{U}=\{u_{1},u_{2},\ldots\}$ connect to the server to send and receive shareable knowledge. Theoretically, TAF can also employ Peer-to-Peer (P2P) decentralized topology if the community is self-organized. In both cases, each participant $u_{k}$ has the exclusive access to its local data $D_{k}$ that is not allowed to be shared for privacy or proprietary reasons. Meanwhile, a certain knowledge exchange protocol has to be established between the participating parties in the first place.
Remark.
A key feature of TAF is that any participant is not necessarily capable of or interested in training. This is because generic knowledge does not always come from model training. For example, participants can swap data statistics, feature rankings, or synthetic samples for their specific purposes.
III-B Formal Definition
The fundamental setting of TAF determines that it is established over “data islands”. This means that any participant $u_{k}$ in the federation $\mathcal{U}$ is associated with a private dataset $D_{k}$. Each participant $u_{k}$ exchanges messages with the coordinator by a customized protocol $\pi_{k}$, which determines what it needs to send and receive and how. The formal definition is given as follows:
Definition 1 (Task-Agnostic Federation (TAF)).
Involving a set of autonomic participants $\mathcal{U}$ each with a private dataset $D_{k}$, TAF is a collaborative system or scenario where each participant $u_{k}$ seeks to maximize its data-related gain $g_{k}(D_{k},\mathcal{U},\pi_{k})$, where $g_{k}$ is a participant-specific function reflecting how much $D_{k}$ is enhanced. Each participant operates by a custom protocol $\pi_{k}$ without knowing any other participant’s objective.
Example. To make the objective more understandable, here we sketch a mock algorithm. First, we use the effective number of samples proposed in [cui2019cbloss] to define $\nu(D)\triangleq(1-\beta^{|D|})/(1-\beta)$, with $\beta$ set as a priori (e.g., $\beta=0.999$). Next, each participant performs data clustering, and reports its centroids to the server in exchange for other participants’ centroids following the protocol $\pi_{k}$ [^4] We are not concerned about privacy leakage in this toy example.. The shared centroids are then used for some local data augmentation $\Theta_{\mathcal{U},\pi_{k}}:D_{k}\rightarrow D^{\prime}_{k}$ (e.g., interpolation), which translates to $u_{k}$’s gain expressed as $g_{k}(D_{k},\mathcal{U},\pi_{k})=\Delta\nu=\nu(D^{\prime}_{k})-\nu(D_{k})$. The gain increases as the algorithm iterates.
Remark.
The gain implies that each participant in TAF is self-motivated and remains agnostic of any task (including its own and other participants’). On this point, TAF differs from pFL and FMTL as they aim to minimize a set of task-specific losses w.r.t. a set of specific models. Further, TAF brings sustainable benefit to the participants since their local data, once enhanced, can serve any downstream purposes beyond training.
Figure 3: (Left) Taxonomy of research related to task-agnostic federation, and (right) the couriers of knowledge employed in the studies.